OpenStudy (anonymous):

HELP MEEEEEE Which equation represents the line passing through the point (4, −16) and is perpendicular to the line

OpenStudy (anonymous):

\[y=-\frac{ 2}{ 3 }x+8\]

OpenStudy (anonymous):

Let us begin with our basic line equation. $$y = mx + b$$

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its already in that form

OpenStudy (anonymous):

Now, we know that, because the line is perpendicular to the first, this line's slope is \(-\frac 1 m\), or the negative reciprocal of that slope.

OpenStudy (anonymous):

So $$y = \frac 3 2 x + b$$ so far.

OpenStudy (anonymous):

We know that x = 4, y = -16 is a solution to this equation, so we can use that to solve for b.

OpenStudy (anonymous):

$$-16 = \frac 3 2 (4) + b$$ What does this simplify to, you think?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Er, the equation simplifies to: $$-16 = \frac{12}2 + b\\ -16 = 6 - b\\ -22 = b$$

OpenStudy (anonymous):

Sorry, that should be 6 + b.

OpenStudy (anonymous):

ooo i thought u was talking about the (4)

OpenStudy (anonymous):

That's one way to do it, but you still have to multiply by 3 in the numerator. Therefore, our equation is: $$y = \frac 3 2 x - 22$$

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

thanx are u up for a few more

OpenStudy (anonymous):

Sure, but I'm not doing all the work this time.

OpenStudy (anonymous):

LMAO OK

OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)?

OpenStudy (anonymous):

we must graph this rite

OpenStudy (anonymous):

Well, you could do that. It might help but it's better to just do it with pure mathematics.